9/11x - 3 = -5/11x - 6
x=99/14
I think I am getting lost in the computation

Answer by fractalier(6550)

Okay, from
(9/11)x - 3 = (-5/11)x - 6
let us add (5/11)x to both sides, then add three to both sides and get
(14/11)x - 3 = -6
(14/11)x = -3
Now multiply both sides by the reciprocal of 14/11, which is 11/14...we get
x = -3(11/14) = -33/14